OpenStudy (anonymous):

(-27)^2/3 write the expression in radical notation B; evaluate the radical expression

OpenStudy (saifoo.khan):

\[\sqrt[3]{-27^2}\] b. 9

jhonyy9 (jhonyy9):

B why is 9 when there is -27 ?

OpenStudy (anonymous):

(-27)^2/3 = \[\sqrt[3]{(-27)^{2}}\] = \[\sqrt[2]{729}\] because (-27) * (-27) = 729

OpenStudy (lalaly):

is it \[\huge{-27^{\frac{2}{3}}}\]or\[\huge{ \frac{-27^2}{3}}\]???

OpenStudy (anonymous):

729 = 9^3 so \[\sqrt[3]{729} = \sqrt[3]{9^{3}} = 9\]